solve the equation │6x +3│= 15
Answer by Edwin McCravy(20056)

            │6x + 3│ = 15

     breaks into two equations:

6x + 3 = 15            6x + 3 = -15
    6x = 12                6x = -18
     x = 2                  x = -3

Two solutions, 2 and -3.

Edwin